OpenStudy (anonymous):

3. 11x + 2y = 65 3x + y = 15 Using the two equations above, solve for x. x = 7 x = -6 x = 30 x = 21

OpenStudy (solomonzelman):

11x + 2y = 65 3x + y = 15 ---> times (-2) --> -6x - 2y = -30 11x + 2y = 65 -6x - 2y = -30 -------------- Add them.